# FuseFlow: A Fusion-Centric Compilation Framework for Sparse Deep Learning on Streaming Dataflow

### v1

@@ Abstract.
As deep learning models scale, sparse computation and specialized dataflow hardware have emerged as powerful solutions to address efficiency. We propose FuseFlow, a compiler that converts sparse machine learning models written in PyTorch to fused sparse dataflow graphs for reconfigurable dataflow architectures (RDAs). FuseFlow is the first compiler to support general cross-expression fusion of sparse operations. In addition to fusion across kernels (expressions), FuseFlow also supports optimizations like parallelization, dataflow ordering, and sparsity blocking. It targets a cycle-accurate dataflow simulator for microarchitectural analysis of fusion strategies. We use FuseFlow for design-space exploration across four real-world machine learning applications with sparsity, showing that full fusion (entire cross-expression fusion across all computation in an end-to-end model) is not always optimal for sparse models—fusion granularity depends on the model itself. FuseFlow also provides a heuristic to identify and prune suboptimal configurations. Using Fuseflow, we achieve performance improvements, including a $\sim$}2.7x speedup over an unfused baseline for GPT-3 with BigBird block-sparse attention.
@@ 1. Introduction
Sparse computation is a promising avenue for greater model efficiency in machine learning (ML) systems (Gale et al., 2020, 2019; Han et al., 2015; Hoefler et al., 2021; Wen et al., 2016), often at the cost of lower hardware utilization. To increase hardware efficiency, researchers and companies are building specialized hardware to accelerate sparse computations (Nowatzki et al., 2017; Carsello et al., 2022; Koul et al., 2024; Dadu et al., 2019; Hegde et al., 2019; Nguyen and Sanchez, 2021; Parashar et al., 2013; Prabhakar et al., 2017; Rucker et al., 2021; Firoozshahian et al., 2023; Siracusa et al., 2023). In order to increase efficiency in these hardware architectures, they are also making increasing use of dataflow, or direct connections between coarse-grained functional units, to rely less on expensive caches, local memories, and memory operations. Dataflow architectures are particularly well-suited for sparse computation because they explicitly coordinate data movement through streaming connections rather than relying on caches, naturally handling the irregular memory access patterns inherent in sparse data (Hsu et al., 2023; Koul et al., 2024; Rucker et al., 2021). Empirically, GPUs are underutilized: a 3-layer GCN inference in PyG on an RTX 5090 across five real-world graphs shows consistently low compute (SM) utilization (avg $16.7\%$) and $\sim$1% memory utilization, with mag as the only case approaching moderate SM utilization (Fig. 1). These observations motivate specialized sparse dataflow accelerators and the compiler support to program them.
@@
Hsu et al. (2023) introduced the Sparse Abstract Machine (SAM) to increase the programmability of these emerging sparse dataflow hardware architectures. The Sparse Abstract Machine is a dataflow abstract machine for sparse tensor algebra computations. It adopts a streaming dataflow model, where data flows between compute nodes. It can express any tensor algebra expression by composing a handful of simple and intuitive dataflow blocks. It also naturally supports expressing fused computation across multiple expressions, different ways to order the dataflow (the iteration order) within an expression (e.g., Gustavson’s algorithm (Gustavson, 1978) versus inner product for sparse matrix multiplication), and lends itself to compile to fabricated hardware (Koul et al., 2024). SAM is therefore a natural starting point as a compiler intermediate representation (IR) for targeting sparse ML models to dataflow hardware.
@@
Hsu et al. (2023) also describe a compilation flow from high-level sparse tensor algebra expressions to SAM dataflow graphs. Their Custard compiler generates SAM graphs that fuse operations across a sparse tensor algebra expression and lets users control the dataflow order. The Custard compilation algorithm is a significant step forward, as the first to demonstrate compilation of sparse tensor algebra expressions to dataflow. It is not, however, suitable for ML model compilation due to its limited capabilities for fusion.
@@
Because Custard compiles individual expressions, it is unable to fuse operations across expressions—a key feature in ML compilers. Moreover, the intra-expression fusion that falls out of Custard’s compilation algorithm fully fuses each expression without any support for partial fusion. Partial fusion is often desirable to provide some of the benefits of fusion while controlling the amount of reuse within a computation (as demonstrated by FlashAttention (Dao et al., 2022)). The tradeoff between fusion and reuse is even more critical, and looks different from that of dense computation, since fused sparse computation may have better asymptotic complexity (Bansal et al., 2023; Hsu et al., 2023; Kjølstad, 2020; Ahrens et al., 2022). Therefore, an ML compiler should expose the fusion granularity as a user schedule so the fusion and reuse tradeoff can be explored across models.
@@
In this paper, we describe a new approach for lowering general sparse ML workloads (beyond pure sparse tensor algebra) to SAM dataflow graphs. Sparse ML refers to deep learning models with one or more sparse tensors. Our approach supports both cross-expression fusion and partial fusion, allowing users to explore the trade-off between fusion and reuse. Our work consists of two new IRs that enable this fusion exploration. The first IR, a fused Einstein Summation (Einsum) representation, tracks the flow of indexing (coordinate) data and values across fused expressions. Then, we introduce a new fusion table representation, a lowering IR that names and memoizes intermediate streams allowing the compiler to reference subgraphs before their materialization to efficiently emit the fused dataflow graph.
@@
We also develop an end-to-end ML compilation system, called FuseFlow, that implements our fusion approach. FuseFlow compiles PyTorch (Ansel et al., 2024) with sparse annotations (Yan et al., 2024; Google, 2021) to SAM graphs. In FuseFlow, users leverage a scheduling language that lets them control fusion granularity and dataflow ordering of expressions. To support modern ML models, we also add support for dense blocks to support block-sparse tensors, non-linear functions, and masking operations to SAM. Finally, our FuseFlow system can generate dataflow graphs that execute in a data-parallel fashion in addition to the pipeline parallelism native to dataflow graphs. Our technical contributions are thus:
@@
  • An algorithm for fusion across multiple independent Einsum expressions (Section 5),
@@
  • A lowering algorithm that converts the fused Einsum expressions to a dataflow representation (Section 6),
@@
We demonstrate the effectiveness of FuseFlow across four model classes by generating 56 equivalent dataflow configurations that yield speedups from ${\sim}1.01$x to ${\sim}3.9$x. Our evaluation underscores the importance of selecting the appropriate fusion granularity and shows that FuseFlow’s heuristic successfully prunes inefficient configurations, offering critical insights for the deployment of large-scale sparse ML applications on dataflow architectures.
@@ 2. Sparse Abstract Machine Background
We provide necessary background on the Sparse Abstract Machine (SAM) (Hsu et al., 2023) to understand our FuseFlow system and the code that it generates. SAM expresses tensor algebra kernels as dataflow graphs by providing a streaming-tensor abstraction and primitives that compose to perform tensor algebra operations. Tensor algebra kernels can be expressed in Einsum notation where tensors are indexed by variables, with addition and multiplication as the core operations. The index variables specify how levels across tensors are broadcast, reduced, and contracted. SAM also introduces the Custard compiler, which compiles high-level Einsum into SAM dataflow graphs. These dataflow graphs are suitable for VLSI implementations and simulation but remain abstract in order to cleanly decouple programs from accelerator implementations.
Figure 2 shows the SAM graph for sparse matrix-vector multiplication (SpMV) $T_{i}=B_{ij}C_{j}$ with the $j\rightarrow i$ dataflow ($\forall_{ji}~T_{i}\mathrel{+}=B_{ij}C_{j}$) where $B$ is stored in a compressed format (e.g., compressed sparse row (CSR)) (Chou et al., 2018; Google, 2021) SAM expresses tensors as streams of data with control tokens, where these tensor streams flow on the arrows between primitives (boxes) in a SAM dataflow graph. SAM’s primitives include:
@@
broadcast operands (e.g., $C$ across each $i$)
@@
elementwise ops and reductions over named indices (e.g., reduce over $j$ in $j\!\rightarrow\!i$).
@@
As in Figure 2, SAM primitives compose together with streams to form SAM graphs that represent any tensor algebra expression with varying dataflows. Arrows in Figure 2 connect dataflow primitives together and transmit streams, where each stream is a sequence of tokens that transmits one level of a tensor in fibertree form (a nested representation of per-level coordinates and values) (Wu et al., 2022). Streams are of three types: coordinates (crd), references to inner levels (ref), and values (val). An $n$-order tensor is represented by $n$ coordinate streams plus one values stream (n+1n{+}1 total).
@@ 3. Forms of Fusion
merges the iteration space of a single tensor algebra expression—co-iterating its inputs—without crossing kernel boundaries. IIF manifests as loop fusion for dense computation, co-iteration for sparse computation, and dataflow iteration fusion in dataflow graphs. Existing sparse tensor compilers that target dataflow hardware (Hsu et al., 2023, 2025) employ this type of fusion, co-iterating to generate fused iteration spaces that elide zeros for one sparse expression at a time.
@@
fuses across different kernels or sub-computations into a single fused computation graph. This type of fusion can be implemented in conjunction with POF and IIF, but is not supported by existing sparse dataflow compilers (Hsu et al., 2023, 2025).
@@
As FuseFlow is a sparse dataflow compiler, we contrast its fusion capabilities with the capabilities of the two prior sparse dataflow compilers Custard (Hsu et al., 2023) and Stardust (Hsu et al., 2025) (C+S) in Figure 4.111The motivation and description in Section 1 is also true for the Stardust compiler (Hsu et al., 2025), another compiler from the same high-level sparse tensor algebra languages to a real dataflow accelerator (Rucker et al., 2021). Custard and Stardust only support fusion within an expression and not across expressions. Although a user can combine expressions into a larger expression, which can then be fused, they must do so by hand and cannot fuse computations that have more than one result. The various fused regions (blue boxes) compare C+S fusion regions with our FuseFlow, which fuses all kernels within a fusion region.
@@
TensorRT (NVIDIA, [n. d.]) POF GPU
XLA (Sabne, 2020) POF, IIF CPU, GPU
DNNFusion (Niu et al., 2021) POF, IIF CPU, GPU
TVM (Chen et al., 2018) POF, IIF CPU, GPU, TPU
TACO (Kjolstad et al., 2017) IIF CPU, GPU
SparseTIR (Ye et al., 2022) IIF CPU, GPU
ReACT (Zhou et al., 2022) IIF CPU, GPU
Stardust (Hsu et al., 2025) IIF Dataflow
Custard (Hsu et al., 2023) IIF Dataflow
@@
In both inter- and intra-expression fusion approaches for sparse computation, there are multiple equivalent ways to iterate through sparse tensors, each with fundamental tradeoffs. Two primary costs, coordinate processing and computation, define the tradeoff between globally fused and factored iteration spaces.
A globally fused iteration space, shown in Figure 5(a), iterates over every index variable, creating an n-dimensional iteration space, where n is the number of index variables (e.g., 4-dimensional in Figure 5(a)). It efficiently filters unnecessary numerical computations but incurs significant coordinate processing overhead, causing coordinate explosion as expressions grow. Prior work on sparse tensor algebra compilation to dataflow accelerators by default generates code that traverses a global iteration space (Hsu et al., 2023, 2025). In contrast, factored iteration iterates pairwise over input tensors (see Figure 5(b)). It generates multiple smaller sub-spaces, one per binary operation (i.e. two 3-dimensional iteration spaces in Figure 5(b)). Each sub-space independently handles coordinate processing, significantly reducing overhead by limiting coordinate analysis to binary operations. However, as this analysis is local rather than global, factored iteration may miss opportunities to skip unnecessary computations, potentially increasing operations.
1for(int i = 0; i < I; i++)
2 for(int k = 0; k < K; k++)
3 for(int j = 0; j < J; j++)
4 for(int l = 0; l < L; l++)
5 D[i,l] += A[i,k] * B[k,j] * C[j,l]
@@ 7 D[i,l] += E[i,j] * C[j,l]
Figure 5. Two iteration patterns for $\forall_{ikjl}C_{il}\mathrel{+}=A_{ik}B_{kj}C_{jl}$, that are represented via loop nests with higher-order reduction variables highlighted in blue (Kjolstad et al., 2019). FuseFlow lowers to a dataflow input iteration graph with a factored iteration space (b), whereas prior work produces dataflow graphs with fully fused iteration spaces (a).

### v2

@@ Abstract.
As deep learning models scale, sparse deep learning (DL) models that exploit sparsity in weights, activations, or inputs and specialized dataflow hardware have emerged as powerful solutions to address efficiency. We propose FuseFlow, a compiler that converts sparse machine learning models written in PyTorch to fused sparse dataflow graphs for reconfigurable dataflow architectures (RDAs). FuseFlow is the first compiler to support general cross-expression fusion of sparse operations. In addition to fusion across kernels (expressions), FuseFlow also supports optimizations like parallelization, dataflow ordering, and sparsity blocking. It targets a cycle-accurate dataflow simulator for microarchitectural analysis of fusion strategies. We use FuseFlow for design-space exploration across four real-world machine learning applications with sparsity, showing that full fusion (entire cross-expression fusion across all computation in an end-to-end model) is not always optimal for sparse models—fusion granularity depends on the model itself. FuseFlow also provides a heuristic to identify and prune suboptimal configurations. Using FuseFlow, we achieve performance improvements, including a $\sim$}2.7x speedup over an unfused baseline for GPT-3 with BigBird block-sparse attention.
sparse machine learning, Einsum, tensor compiler, kernel fusion, spatial dataflow accelerator
journalyear: 2026copyright: ccconference: Proceedings of the 30th ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 2; March 21–26, 2026; Pittsburgh, PA, USA.booktitle: Proceedings of the 30th ACM International Conference on Architectural Support for Programming Languages and Operating Systems, Volume 2 (ASPLOS ’26), March 21–26, 2026, Pittsburgh, PA, USAisbn: 979-8-4007-2359-9/2026/03doi: 10.1145/3779212.3790165ccs: Computer systems organization Data flow architecturesccs: Theory of computation Streaming modelsccs: Computing methodologies Machine learningccs: Software and its engineering Compilers
@@ 1. Introduction
Deep learning models may have sparse weights, activations, or inputs—naturally occurring or induced. Exploiting this sparsity during computation, which we call sparse deep learning (DL), reduces compute and memory requirements but introduces irregular memory access patterns (gale2020; gale2019state; han2015learningweightsconnectionsefficient; hoefler2021sparsity; wen2016learning). To increase hardware efficiency, researchers are building specialized hardware to accelerate sparse computations (nowatzki2017; carsello2022amber; onyx; dadu2019spu; hegde2019extensor; fifer2021nguyen; triginst2013parashar; plasticine; rucker2021capstan; firoozshahian2023mtia; siracusa2023tensor). In order to increase efficiency in these hardware architectures, they are also making increasing use of dataflow, or direct connections between coarse-grained functional units, to rely less on expensive caches, local memories, and memory operations. Dataflow architectures are particularly well-suited for sparse computation because they explicitly coordinate data movement through streaming connections rather than relying on caches, naturally handling the irregular memory access patterns inherent in sparse data (hsu2023sam; onyx; rucker2021capstan). Empirically, GPUs are underutilized: A 3-layer GCN inference in PyTorch Geometric (PyG) on an RTX 5090 across five real-world graphs shows consistently low compute (SM) utilization (avg $16.7\%$) and $\sim$1% memory utilization (Figure 1). These observations motivate specialized sparse dataflow accelerators and the compiler support to program them.
@@
hsu2023sam introduced the Sparse Abstract Machine (SAM) to increase the programmability of these emerging sparse dataflow hardware architectures. The Sparse Abstract Machine is a dataflow abstract machine for sparse tensor algebra computations. It adopts a streaming dataflow model, where data flows between compute nodes. It can express any tensor algebra expression by composing a handful of simple and intuitive dataflow blocks. It also naturally supports expressing fused computation across multiple expressions, different ways to order the dataflow (the iteration order) within an expression (e.g., Gustavson’s algorithm (gustavson1978) versus inner product for sparse matrix multiplication), and lends itself to compile to fabricated hardware (onyx). SAM is therefore a natural starting point as a compiler intermediate representation (IR) for targeting sparse ML models to dataflow hardware.
@@
hsu2023sam also describe a compilation flow from high-level sparse tensor algebra expressions to SAM dataflow graphs. Their Custard compiler generates SAM graphs that fuse operations across a sparse tensor algebra expression and lets users control the dataflow order. The Custard compilation algorithm is a significant step forward, as the first to demonstrate compilation of sparse tensor algebra expressions to dataflow. It is not, however, suitable for ML model compilation due to its limited capabilities for fusion.
@@
Because Custard compiles individual expressions, it is unable to fuse operations across expressions—a key feature in ML compilers (lacouturechallenges; Chen2018tvm). Moreover, the intra-expression fusion that falls out of Custard’s compilation algorithm fully fuses each expression without any support for partial fusion. Partial fusion is often desirable to provide some of the benefits of fusion while controlling the amount of reuse within a computation (as demonstrated by FlashAttention (dao2022flashattention)). The fusion–recomputation tradeoff is fundamental: fusion eliminates intermediate tensors between operations, reducing memory traffic, but excessive fusion can force recomputation of values (zhou2022react; dias2022sparselnr). Conversely, insufficient fusion materializes intermediate tensors to memory, forcing more data movement. This tradeoff is even more critical, and looks different from that of dense computation, since fused sparse computation may have better asymptotic complexity (bansal2023mosaic; hsu2023sam; kjolstad2020sparse; ahrens2022autoscheduling). Therefore, an ML compiler should expose the fusion granularity as a user schedule so the fusion and reuse tradeoff can be explored across models.
In this paper, we describe a new approach for lowering sparse DL models, models with one or more sparse tensors, to SAM dataflow graphs. Unlike prior work on sparse tensor algebra compilers that target individual expressions, FuseFlow compiles complete sparse DL inference pipelines, including nonlinear operations and masking. FuseFlow supports sparse tensors from any source, whether from pruning, natural zeros, or induced patterns, provided that the sparse data structure type is determined before compilation (Section 4.1).
@@
Our approach supports both cross-expression fusion and partial fusion, allowing users to explore the trade-off between fusion and reuse. Our work consists of two new IRs that enable this fusion exploration. The first IR, a fused Einstein Summation (Einsum) representation, tracks the flow of indexing (coordinate) data and values across fused expressions. Then, we introduce a new fusion table representation, a lowering IR that names and memoizes intermediate streams allowing the compiler to reference subgraphs before their materialization to efficiently emit the fused dataflow graph.
@@
We also develop FuseFlow, the first academic end-to-end sparse ML compiler for reconfigurable dataflow architectures. FuseFlow compiles PyTorch (pytorch2) with sparse annotations (yan2024scorch; mpact) to SAM graphs. In FuseFlow, users leverage a scheduling language that lets them control fusion granularity and dataflow ordering of expressions. To support modern ML models, we also add support for dense blocks to support block-sparse tensors, non-linear functions, and masking operations to SAM. Finally, our FuseFlow system can generate dataflow graphs that execute in a data-parallel fashion in addition to the pipeline parallelism native to dataflow graphs. FuseFlow targets both cycle-accurate simulation and FPGA synthesis and existing dataflow accelerators (onyx; chen2025opal), with validated agreement between FPGA and the simulation (Section 8.2). Our technical contributions are thus:
@@
  • A new data structure and algorithm for fusion across multiple independent Einsum expressions (Section 5),
  • A new abstraction that enables interleaved reductions for factored iteration and on-the-fly rearrangement of dataflow graphs (Section 6),
@@
  • A lowering algorithm that converts the fused Einsum expressions to a dataflow representation (Section 6)
@@
We demonstrate the effectiveness of FuseFlow across four model classes by generating 56 equivalent dataflow configurations that yield speedups from ${\sim}1.5$x to ${\sim}3.9$x. Our evaluation underscores the importance of selecting the appropriate fusion granularity and shows that FuseFlow’s heuristic successfully prunes inefficient configurations, offering critical insights for the deployment of large-scale sparse ML applications on dataflow architectures.
@@ 2. Sparse Abstract Machine Background
We provide necessary background on the Sparse Abstract Machine (SAM) (hsu2023sam) to understand our FuseFlow system and the code that it generates. SAM expresses tensor algebra kernels as dataflow graphs by providing a streaming-tensor abstraction and primitives that compose to perform tensor algebra operations. Tensor algebra kernels can be expressed in Einsum notation where tensors are indexed by variables, with addition and multiplication as the core operations. The index variables specify how levels across tensors are broadcast, reduced, and contracted. SAM also introduces the Custard compiler, which compiles high-level Einsum into SAM dataflow graphs. These dataflow graphs are suitable for VLSI implementations and simulation but remain abstract in order to cleanly decouple programs from accelerator implementations.
Figure 2 shows the SAM graph for sparse matrix-vector multiplication (SpMV) $T_{i}=B_{ij}C_{j}$ with the $j\rightarrow i$ dataflow ($\forall_{ji}~T_{i}\mathrel{+}=B_{ij}C_{j}$) where $B$ is stored in a compressed format (e.g., compressed sparse row (CSR)) (chou2018; mpact) SAM expresses tensors as streams of data with control tokens, where these tensor streams flow on the arrows between primitives (boxes) in a SAM dataflow graph. SAM’s primitives include:
@@
broadcast operands (e.g., $C$ across each $i$).
@@
perform elementwise operations and reductions (e.g., reduce over $j$ in $j\!\rightarrow\!i$).
@@
As in Figure 2, SAM primitives compose together with streams to form SAM graphs that represent any tensor algebra expression with varying dataflows. Arrows in Figure 2 connect dataflow primitives together and transmit streams, where each stream is a sequence of tokens that transmits one level of a tensor in fibertree form (a nested representation of per-level coordinates and values) (sparseloop). Streams are of three types: coordinates (crd), references to inner levels (ref), and values (val). An $n$-order tensor is represented by $n$ coordinate streams plus one values stream (n+1n{+}1 total).
@@ 3. Forms of Fusion
merges the iteration space of a single tensor algebra expression—co-iterating its inputs—without crossing kernel boundaries. IIF manifests as loop fusion for dense computation, co-iteration for sparse computation, and dataflow iteration fusion in dataflow graphs. Existing sparse tensor compilers that target dataflow hardware (hsu2023sam; hsu2025stardust) employ this type of fusion, co-iterating to generate fused iteration spaces that elide zeros for one sparse expression at a time.
@@
fuses across different kernels or sub-computations into a single fused computation graph. This type of fusion can be implemented in conjunction with POF and IIF, but is not supported by existing sparse dataflow compilers (hsu2023sam; hsu2025stardust).
@@
As FuseFlow is a sparse dataflow compiler, we contrast its fusion capabilities with the capabilities of the two prior sparse dataflow compilers Custard (hsu2023sam) and Stardust (hsu2025stardust) (C+S) in Figure 4.111The motivation and description in Section 1 is also true for the Stardust compiler (hsu2025stardust), another compiler from the same high-level sparse tensor algebra languages to a real dataflow accelerator (rucker2021capstan). Custard and Stardust only support fusion within an expression and not across expressions. Although a user can combine expressions into a larger expression, which can then be fused, they must do so by hand and cannot fuse computations that have more than one result. The various fused regions (blue boxes) compare C+S fusion regions with our FuseFlow, which fuses all kernels within a fusion region.
@@
TensorRT (tensorrt) POF GPU
XLA (sabne2020xla) POF, IIF CPU, GPU
DNNFusion (niu2021dnnfusion) POF, IIF CPU, GPU
TVM (Chen2018tvm) POF, IIF CPU, GPU, TPU
TACO (kjolstad2017taco) IIF CPU, GPU
SparseTIR (sparsetir) IIF CPU, GPU
ReACT (zhou2022react) IIF CPU, GPU
Stardust (hsu2025stardust) IIF Dataflow
Custard (hsu2023sam) IIF Dataflow
@@
In both inter- and intra-expression fusion approaches for sparse tensor operations, there are multiple equivalent ways to iterate through sparse tensors, each with fundamental tradeoffs. Two primary costs, coordinate processing and computation, define the tradeoff between globally fused and factored iteration spaces.
A globally fused iteration space, shown in Figure 5(a), iterates over every index variable, creating an n-dimensional iteration space, where n is the number of index variables (e.g., 4-dimensional in Figure 5(a)). It efficiently filters unnecessary numerical computations but incurs significant coordinate processing overhead, causing coordinate explosion as expressions grow. Prior work on sparse tensor algebra compilation to dataflow accelerators by default generates code that traverses a global iteration space (hsu2023sam; hsu2025stardust). In contrast, factored iteration iterates pairwise over input tensors (see Figure 5(b)). It generates multiple smaller sub-spaces, one per binary operation (i.e. two 3-dimensional iteration spaces in Figure 5(b)). Each sub-space independently handles coordinate processing, significantly reducing overhead by limiting coordinate analysis to binary operations. However, as this analysis is local rather than global, factored iteration may miss opportunities to skip unnecessary computations, potentially increasing operations.
for(int i = 0; i < I; i++)
[…]
for(int l = 0; l < L; l++)
D[i,l] += A[i,k] * B[k,j] * C[j,l]
@@
Figure 5. Two iteration patterns for $\forall_{ikjl}C_{il}\mathrel{+}=A_{ik}B_{kj}C_{jl}$, that are represented via loop nests with higher-order reduction variables highlighted in blue (kjolstad2019workspaces). FuseFlow lowers to a dataflow input iteration graph with a factored iteration space (b), whereas prior work produces dataflow graphs with fully fused iteration spaces (a).